The length of a rectangle is 23 centimeters less than six times its width. Its area is 35 square centimeters. Find the dimensions of the rectangle.
Answer by Boreal(15235)

width is x cm
length is 6x-23 cm
area is 35 cm^2
6x^2-23x=35
6x^2-23x-35=0
x^2-23x-210=0
(x-30)(x+7)=0
divide by 6 and reduce
(x-5)(x+7/6)=0
(x-5)(6x+7)=0
only positive root is x=5 cm
6x-23=30-23=7 cm LENGTH
5 x 7 cm